A builder says he can construct an office building for about $65 a square foot. What is the estimated cost for a 1-story building that is 50 feet wide and 67 feet long?
1
Expert's answer
2011-06-17T07:49:06-0400
The floor space is S = 50*67 = 3350 ft2,
The estimated cost would be:
C = $65 * 3350 = $217750.